Mr. Jackons is 13 years less than 3 times as old as his son. six years ago he was 14 years more than twice as old as his son was then. Find thire age
Answer by Paul(988)

3x-13=-6+14+2x
3x-13=8+2x
3x-13+13=8+13+2x
3x=21+2x
3x-2x=21+2x-2x
x=21
3(21)-13=50
Hence his son is 21 and he is 50